(1 point) Book Problem 27 Consider the function f(x) = 12x5 + 15x4 - 240x3 + 6. This function has 3 critical numbers A <B<C:Each blank is a part of one question so fill in all blanks please.
Answer #1


fla)= 1245+ 15x4 - 240x3 +6 f(x)= 60x4 + 6003 - 720 x2 For Critical points flra)=0 T 60x4 +603-720 x2=0 60x2 x2 + 4x -3x-12} This function has 3 critical numbers ACB cc: 3 c - A = -4 B = Now local Minima + 3 local Maxina -4 o Trotterties print (Neithfla) has inflection points at x=D, EIF D = 3-5393 F = -3 + 393 8 8 Now , + + F D E on (-os ID) TCD [DE] CU CE,P] : C D n CU [